OpenStudy (anonymous):

How do you simplifiy the expression: (1+ cot theta)(1-cot theta) - csc^2 theta ?

OpenStudy (anonymous):

could give clear eqautin

OpenStudy (mathteacher1729):

Perfect square binomial: \[(1-x)(1+x)=1-x^2\] In this case "x" = \(\cot(\theta)\). So that means \( (1+\cot (\theta))(1-\cot (\theta)) = 1 - \cot^2(\theta)\) The expression is now \(1-\cot^2(\theta) - \csc^2(\theta)\). We can use some trig identities here to reduce \(1-csc^2(\theta) = -cot^2(\theta)\) Now our final answer is almost in front of us. :)

OpenStudy (anonymous):

what math teacher said. nice

OpenStudy (anonymous):

u reduce it further by simplifying csc and cot..?

OpenStudy (anonymous):

i completely understand the first part! thanks!

OpenStudy (anonymous):

I'm just saying..its not an equation. it needs to be simplified to an expression

OpenStudy (anonymous):

rewrite everything (after last step) in terms of sine and cosine and it will be easier

OpenStudy (anonymous):

okay I will. thanks!

OpenStudy (anonymous):

so the answer is (1 - cot^2 theta) - csc theta right?

OpenStudy (anonymous):

nevermind. 2cot^2 theta right?

OpenStudy (anonymous):

well i guess there are many ways to write this. i got

OpenStudy (anonymous):

we start with \[-cot^2(\theta)-csc^2(\theta)\]

OpenStudy (anonymous):

wait can I guess. thank you btw! i really appreciate your help. -2cot^2 theta

OpenStudy (anonymous):

hold on i forgot the original question

OpenStudy (anonymous):

yes! that is what i got after confusing myself. good work

OpenStudy (anonymous):

ahhh thank you so much!!!!!